OpenStudy (anonymous):

The electrical resistance R of a wire varies directly as the length I of the wire and inversely as the square of its diameter d. If 20 meters of wire of diameter 1.5 mm has a resistance of 12 ohms, what is the resistance of 20 meters of the same wire if the diameter is increased to 5 mm?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

R= k L/d^2 k 20/(1.5)^2=12 let's find k k=1.35 R=1.35 L/D^2 = 1.35 (20/25)

OpenStudy (anonymous):

R=1.08

OpenStudy (anonymous):

Thank you so much :)